OpenStudy (anonymous):

what is 2(9n-1)+7(9n+6)=35

OpenStudy (anonymous):

2(9n-1)+7(9n+6)=35<----- Using the distribution property... 18n - 2 + 63n + 42 = 35<----- Commutative property now... 81n + 40 = 35 - 40 - 40 ---------------- 81n = -5 /81 /81 -------------- n = -5/81

OpenStudy (anonymous):

First expand the brackets

OpenStudy (anonymous):

Distribute the 2 and the 7 through the parenthesis and then rearrange the equation such that you have the numbers that have n attached to them on one side of the equality sign and the rest on the other side. Simplify that first and then isolate n to find the answer

OpenStudy (anonymous):

And group like terms

OpenStudy (anonymous):